### v1

# MSI-NeRF: Linking Omni-Depth with View Synthesis through Multi-Sphere Image aided Generalizable Neural Radiance Field

Abstract
Panoramic observation using fisheye cameras is significant in robot perception, reconstruction, and remote operation. However, panoramic images synthesized by traditional methods lack depth information and can only provide three degrees-of-freedom (3DoF) rotation rendering in virtual reality applications. To fully preserve and exploit the parallax information within the original fisheye cameras, we introduce MSI-NeRF, which combines deep learning omnidirectional depth estimation and novel view rendering. We first construct a multi-sphere image as a cost volume through feature extraction and warping of the input images. It is then processed by geometry and appearance decoders, respectively. Unlike methods that regress depth maps directly, we further build an implicit radiance field using spatial points and interpolated 3D feature vectors as input. In this way, we can simultaneously realize omnidirectional depth estimation and 6DoF view synthesis. Our method is trained in a semi-self-supervised manner. It does not require target view images and only uses depth data for supervision. Our network has the generalization ability to reconstruct unknown scenes efficiently using only four images. Experimental results show that our method outperforms existing methods in depth estimation and novel view synthesis tasks.
I Introduction
Omnidirectional imagery is an important tool for robotics, computer vision, and virtual reality (VR). Compared to normal pin-hole images with limited field-of-view (FoV), omnidirectional images include all viewing angles both horizontally and vertically. It creates a compact way to store information about an entire scene in just one camera shot. Traditionally, omnidirectional images are synthesized using multiple fisheye cameras or camera arrays [1, 2]. Seamless omnidirectional images can be generated by matching and stitching overlapping pixels between adjacent images. They are then saved in equirectangular projected images using spherical coordinate representation.
Although the abovementioned method can synthesize omnidirectional images with pleasant visual quality, more is needed. On the one hand, omnidirectional images assume that the rays corresponding to all pixels have the same origin. Since it is impossible to place the optical centers of multiple cameras at the same point, there will always be errors in the stitching method, especially in large parallax scenarios. On the other hand, stitching and blending images from cameras at different positions will cause the originally existing disparity information to be eliminated. As a result, the final produced image loses its depth information and can only be rendered in 3DoF, limiting its practical usage in the scope of robotics and virtual reality.
[…]
With the development of computer vision and deep learning, some methods try to solve such problems. OmniMVS [3] extends the traditional multi-view stereo (MVS) method to panoramic views. It does not directly stitch images together but uses deep learning methods to match features between different cameras and returns the panoramic depth map. However, it does not consider occluded regions and is insufficient to support 6DoF rendering. OmniNeRF [4] proposes a method for 6DoF rendering panoramically. However, since it requires processed omnidirectional images and additional depth as input, it can not be implemented end-to-end. The need for per-scene optimization also limits its application in fields with high real-time requirements. MVSNeRF [5] achieves free view rendering with generalizable network. Nevertheless, it only works for object-level or front-facing scene reconstruction. No research still achieves the reconstruction of panoramic scenes using a single shot.
This paper aims to generate an omnidirectional representation that preserves the 3D information in the input multi-view fisheye images. We can obtain omnidirectional depth information from such representation and perform 6DoF perspective synthesis. As input, we collect images from four surrounding fisheye cameras. Then, we follow the general MVS pipeline to build a multi-sphere image (MSI) at predefined depth layers as a cost volume. A NeRF [6] is then constructed to represent the scene with an implicit function. It takes the interpolated features from the cost volume as an additional input to better utilize the learned geometry and texture. For network training, our method utilizes semi-self-supervision. The traditional training pipeline cannot meet the needs since mainstream panoramic datasets only contain depth ground truth and do not include target views for training. Our method utilizes input images to introduce color supervision into the network, enabling generalizable NeRF training using only depth ground truth. The outline of our method is shown in Fig. 1.
With our approach, we can efficiently render omnidirectional depth maps and novel view images. Vision-based panoramic depth has the advantages of lower cost, higher resolution, and larger FoV compared with LiDAR. Such depth data plays an important role in autonomous driving, robot perception, and scene reconstruction. For the field of virtual reality, being able to render 6DoF novel views eliminates the VR sickness caused by 3DoF rendering. It also benefits the development of human-machine remote collaboration and teleoperation.
In summary, our main contributions are as follows:
  • $\bullet$
    We propose a deep learning method to synthesize an omnidirectional radiance field from only four fisheye inputs. The traditional 2D panorama output is expanded into 3D, and the parallax information inside the original images is preserved.
  • $\bullet$
    We combine the depth estimation task with the novel view synthesis task. By building a semi-self-supervised framework, multi-task network training can be achieved only using the most common depth data supervision.
  • $\bullet$
    Our trained network can generalize across scenes and achieve an inference speed of approximately ten frames per second. In experiments, our method performs well on both tasks and achieves state-of-the-art results.
II Related Works
II-A Omnidirectional Depth Estimation
Depth estimation using multi-view images has long been a concern in computer vision and robotics. By stereo-matching colors or features from different viewing angles, depth information can be obtained through the MVS methods [7, 8].
Unlike the general MVS algorithms, which can only generate depth maps of a specific target perspective, omnidirectional depth estimation aims to obtain panoramic depth using cameras surrounding all $360^{\circ}$. It usually requires multiple large FoV fisheye cameras [9, 3, 10, 11] or spherical camera array [12, 13] as input. SweepNet [9] modifies the Plane Sweep method [11] into spherical coordinates and uses Semi-Global Matching (SGM) [14] to obtain a depth map. Its following work, OmniMVS [3], also uses the quad fisheye array but improves the matching module to a 3D convolution network. Sphere-Sweeping-Stereo [10] proposes to use adaptive spherical matching and cost aggregation to accelerate inference while maintaining a fine-grained depth map quality.
The above omnidirectional depth estimation methods can, to some extent, solve the problems of parallax information utilization and 3D scene perception. However, free view roaming and rendering occluded parts can still not be achieved.
II-B Omnidirectional Novel View Synthesis
Novel view synthesis means to continuously interpolate new views given a limited set of source views. NeRF [6] and multi-plane image (MPI) [15] methods perform view synthesis on object level and forward facing scenes by constructing implicit or explicit geometric proxy. Although high quality novel view images can be generated, the bounded scene region still limits their practical usage.
In order to extend novel view synthesis to unbounded scenes, methods using full space mapping [16, 17] and omnidirectional input [4, 18] have been developed. Mip-NeRF 360 [16] enhances Mip-NeRF [19] with scene parameterization to realize $360^{\circ}$ scene representation. But it still needs densely captured views for high quality rendering. Methods taking omnidirectional images alleviates such problem. MatryODShka [18] takes omnidirectional stereo (ODS) images as input and uses neural network to blend them into MSI for novel view synthesis. OmniNeRF [4] first synthesizes multi-view images from color and depth panorama by point projection. It then trains a scene specific NeRF using the two-stage method, which lacks practicality.
Due to that multi-view datasets containing panoramic images are difficult to obtain, methods usually use single view with depth auxiliary for nerf training. However, these methods usually lead to view penetration and artifacts.
Figure 2: Structure of our method. Our method can be divided into three parts. First, by 2D feature extraction and warping, a multi-sphere image (MSI) representation can be built. Then, through geometric and appearance 3D decoder, explicit features ($\mathbf{f}_{geo}$, $\mathbf{f}_{appr}$) containing prior information can be obtained. They are then fed into NeRF implicit MLP along with point position $\mathbf{x}$, ray direction $\mathbf{d}$, and projected color $\mathbf{c}_{proj}$. The output occupancy and color are used for fisheye color image and depth image rendering. The learnable parameters in CNN and MLP are trained through semi-self-supervision.
II-C Generalizable NeRF
NeRF and its derived algorithms have shown great performance in 3D scene representation. However, most of them require scene-specific training. Whenever faced with a new scenario, these methods need seconds or even hours of training before they can generate high-quality results [20]. The nature of training from scratch also makes them dependent on dense source views to ensure that novel views can be correctly interpolated.
Leveraging the development of deep learning and stereo matching, NeRF with generalizability gradually emerged [21, 22, 5]. These generalizable NeRFs eliminate scene-specific training and can quickly synthesize novel views after training on large-scale datasets, even with only sparse view inputs. PixelNeRF [21] first uses a CNN encoder to extract image features from source views. It then builds an implicit function that takes the projected 2D features as additional input, following the idea of image-based rendering. IBRNet [22] proposes a similar way to achieve generalizability. It uses a ray transformer to add interaction along the ray during the render process. Differently, MVSNeRF [5] incorporates MVS methods into NeRF rendering. It first constructs a cost volume from the source views using plane sweeping and 3D convolution. Interpolated 3D features and the projected colors are then added to the NeRF input, transferring prior knowledge into the implicit field.
Although the above methods achieve fast generalizable NeRF reconstruction using only a small number of views, they are all targeted at the object level or forward-facing scene situations. Obtaining omnidirectional scene representation using a single shot remains a challenging problem.
III Method
In this work, we present MSI-NeRF, a deep learning method that inputs four fisheye images and outputs a panoramic scene representation. It adopts the representation form of NeRF and uses priorly generated multi-sphere images as a geometric aid. Such representation enables 6DoF view synthesis while retaining depth information. During optimization, our method only requires depth ground truth. It can introduce color information into the network through input data to achieve semi-self-supervision. Our MSI-NeRF consists of three blocks: multi-sphere image construction (Sec. III-A), hybrid neural rendering (Sec. III-B), and multi-task semi-self-supervision (Sec. III-C). Its overall structure is shown in Fig. 2.
III-A Multi-Sphere Image Construction
Our method takes four wide FoV fisheye images as input. They are captured from a multi-fisheye camera rig with four cameras surrounding the $360^{\circ}$ of the scene. First, a 2D CNN network with sharing weights extracts feature maps from the captured images. Then, a multi-sphere image is built to aggregate the multi-view images. It consists of a series of predefined depth layers. Each layer is parameterized by equirectangular projection with the same resolution. In order to handle scenes with infinite depth while considering the reconstruction precision of close-range scenes, we uniformly divide these layers by their inverse depth. Similar to the plane sweeping methods done in the forward-facing scenes, we project pixels on the MSI back to every fisheye camera using their intrinsics and extrinsics. The value of each pixel will be filled with the concatenation of the feature vectors from the four feature maps. The projection is shown in the following equations:
$\overline{\mathbf{p}}(\theta,\phi)=(\cos(\phi)\cos(\theta),\sin(\phi),\cos(\phi)\sin(\theta))^{\top},$ (1)
$\mathbf{V}_{n,i}(\theta,\phi)=F_{i}\left(\Pi_{i}\left(\overline{\mathbf{p}}(\theta,\phi)/d^{-1}_{n}\right)\right),$ (2)
where $\overline{\mathbf{p}}$ represents the unit ray for the spherical coordinate $<\theta,\phi>$ of the equirectangular projection, $\Pi_{i}$ is the projection function from rig coordinate to the $i^{th}$ camera, and $d^{-1}_{n}$ is the inverse depth of the $n^{th}$ layer. $\mathbf{V}$ alias the generated spherical volume representation, and $F$ is the image feature map.
Finally, the constructed spherical layers $\mathbf{V}$ will be processed by 3D CNN networks to match and fuse the features from different viewing angles. Unlike OmniMVS [3], which directly regress the cost volume in the depth dimension to obtain a depth map, and MatryODShka [18], which directly blends colors between these layers, we aim to build a continuous 3D representation that contains both geometry and appearance information. Specifically, we use two separate networks to further process the obtained MSI volume $\mathbf{V}$. The first network works as a comparator, generating a cost volume to store geometric information. The second network additionally outputs an appearance volume, retaining the projected color information. Their functionality can be demonstrated as the following process:
$\mathbf{V}_{geo}=h_{geo}(\mathbf{V}),\quad\mathbf{V}_{app}=h_{app}(\mathbf{V}),$ (3)
where $h_{geo}$ and $h_{app}$ represents the two decoders, and $\mathbf{V}_{geo}$ and $\mathbf{V}_{app}$ are the generated feature volumes. By aggregating the feature vectors from different views and conducting information interaction with neighboring spatial voxels, reasonable space occupancy, and color can be encoded in the resulting 3D feature grids $\mathbf{V}_{geo}$ and $\mathbf{V}_{app}$. They will be used as explicit aids for radiance field construction and neural rendering. By distinguishing between appearance and geometry volume representations, we can make our decoder network functions more concrete, each performing their own duties. It also allows interaction and clarity of network learning.
III-B Hybrid Neural Rendering
In order to achieve free viewpoint roaming, continuous representation is needed to infer color and geometric attributes at any location. We utilize the cutting-edge NeRF [6] for viewpoint interpolation and rendering. However, vanilla NeRF lacks the generalization ability and can only preserve knowledge of a single scene. In this work, we incorporate generalizable knowledge from front-end deep neural networks trained on large-scale datasets with NeRF implicit function and form a hybrid neural rendering system.
We construct our implicit function using an MLP. It can be represented as:
$o,\mathbf{c}=f(\mathbf{x},\mathbf{d},\mathbf{f}_{geo},\mathbf{f}_{appr},\mathbf{c}_{proj},\Theta),$ (4)
where $o$ and $\mathbf{c}$ are the outputted occupancy probability and RGB color, $\Theta$ is the network parameter that implicitly saves the scene information, and $\mathbf{x}$ and $\mathbf{d}$ represent the spatial coordinate and viewing direction respectively. The interpolated geometry and appearance features $\mathbf{f}_{geo}$ and $\mathbf{f}_{appr}$ carrying prior learned information from the MSI volume $\mathbf{V}_{geo}$ and $\mathbf{V}_{app}$ are also fed into the function. We additionally consider the projected pixel colors $\mathbf{c}_{proj}$ from the original fisheye images as input, providing a hint for color learning. It can also be viewed as a way of implicit color blending. The above variables are obtained by the following equation:
$\mathbf{f}_{geo},\mathbf{f}_{appr}$ $=S(\mathbf{x},\mathbf{V}_{geo},\mathbf{V}_{app}),$ (5)
$\mathbf{c}_{proj,i}$ $=I_{i}\left(\Pi_{i}\left(\mathbf{x}\right)\right),$
where $S$ represents the interpolation operation. In particular, we first apply positional encoding on the position and direction vectors. Then, we use an MLP taking embedded $\mathbf{x}$, $\mathbf{d}$, and $\mathbf{f}_{geo}$ as input, generating the occupancy and a feature vector. Finally, the feature vector is then processed by another MLP along with $\mathbf{f}_{appr}$ and $\mathbf{c}_{proj}$, generating the final color output.
Since we use inverse depth as the MSI layer arrangement criterion, for distant layers, the distance between two adjacent layers is also greater. In this way, using 3D interpolation will cause wrongly mixed prior features and affect convergence during training. As a result, we use the intersection of rays and spherical depth layers as our sample points. This way, only 2D interpolation on the 3D volume is needed, avoiding blending between depth layers with large gaps while maintaining balanced sample points. The intersection is calculated as follows:
$a=\mathbf{r}_{d}^{T}\mathbf{r}_{d},\quad b=2\mathbf{r}_{d}^{T}\mathbf{r}_{o},\quad c_{n}=\mathbf{r}_{o}^{T}\mathbf{r}_{o}-1/(d^{-1}_{n})^{2},$ (6)
$z_{n}=\frac{-b+\sqrt{b^{2}-4ac_{n}}}{2a}\quad\text{if}\quad\left\|\mathbf{r}_{o}\right\|_{2}<1/(d^{-1}_{n}),$ (7)
where $\mathbf{r}_{o}$ and $\mathbf{r}_{d}$ are the origin and unit direction of the ray $\mathbf{r}$, and $z_{n}$ is the depth of the intersecting point with the $n^{th}$ sphere along the ray. Note that the equation only applies to the case where rays are emitted inside the sphere. After points sampling, we follow the pipeline of UNISURF [23] to conduct volume rendering:
$\hat{\mathbf{C}}(\mathbf{r})=\sum_{i=1}^{N_{s}}w\left(\mathbf{p}_{i}\right)\mathbf{c}\left(\mathbf{p}_{i},\mathbf{r}\right),\quad\hat{D}(\mathbf{r})=\sum_{i=1}^{N_{s}}w\left(\mathbf{p}_{i}\right)d\left(\mathbf{p}_{i}\right),$ (8)
where $N_{s}$ is the number of intersecting spheres, $\mathbf{p}_{i}$ denotes the $i^{th}$ sampled points on ray $\mathbf{r}$, and $d(\mathbf{p}_{i})$ denotes the depth value at position $\mathbf{p}_{i}$. The weight $w\left(\mathbf{p}_{i}\right)$ is the product of the occupancy value $o(\mathbf{p}_{i})$ and the occlusion term $T\left(\mathbf{p}_{i}\right)$:
$w\left(\mathbf{x}_{i}\right)=o\left(\mathbf{p}_{i}\right)T\left(\mathbf{p}_{i}\right),\quad T\left(\mathbf{p}_{i}\right)=\prod_{j=1}^{i-1}\left(1-o\left(\mathbf{p}_{j}\right)\right).$ (9)
Through the volume rendering, we can synthesize color images $\hat{\mathbf{C}}(\mathbf{r})$ and depth maps $\hat{D}(\mathbf{r})$ at any viewpoint with any camera intrinsics.
III-C Multi-Task Semi-Self-Supervision
In order to let our network generalize across different scenes, training on large-scale panoramic image datasets is needed. However, large-scale datasets containing source fisheye images as input and multi-view target images as ground truth are difficult to obtain. Such difficulties will lead to trouble in learning scene appearance and novel view synthesis. In this work, we utilize the widely used fisheye omnidirectional depth estimation datasets proposed by OmniMVS [3] for both geometry and appearance understanding of panoramic scenes. We revisit the training mechanism of generalizable NeRFs and find that the purpose of using multi-view target images for network training is: (a) to predict scene occupy information and (b) to predict the appearance information of the occluded part of the scene. These two points allow generalizable NeRFs to quickly synthesize novel views from a small number of target views, using the learned prior knowledge. Now that we already have the depth supervision to learn occupancy, we only need to find a way to inject occluded color supervision into training. We come up with the idea of directly using the input fisheye images as a multi-view supervision to provide texture information for the occluded parts.
To this end, we propose our multi-task semi-self-supervision strategy. For the original depth ground truth aligned with the rig coordinate, we use the rendered depth to compare it as depth supervision. We further render fisheye color images using the source camera’s intrinsic and extrinsic parameters and compare them with the captured images, forming an additional color supervision. Our loss function is represented as:
$\mathcal{L}_{depth}$ $=\frac{1}{|\mathcal{R}_{p}|}\sum_{\mathbf{r}\in\mathcal{R}_{p}}\left\|\hat{D}(\mathbf{r})-D(\mathbf{r})\right\|_{1},$ (10)
$\mathcal{L}_{color}$ $=\frac{1}{|\mathcal{R}_{f}|}\sum_{\mathbf{r}\in\mathcal{R}_{f}}\left\|\hat{\mathbf{C}}(\mathbf{r})-\mathbf{C}(\mathbf{r})\right\|_{2}^{2},$
$\mathcal{L}=\mathcal{L}_{color}+\lambda_{d}\mathcal{L}_{depth},$ (11)
where $\mathcal{R}_{p}$ is the set of rays that makes up the central panoramic depth image, and $\mathcal{R}_{f}$ is the set of rays that constitutes the four fisheye images.
If only input fisheye images are used for supervision, the network will suffer from serious over-fitting and lose the ability to predict scene information from other perspectives. However, after adding depth supervision, correct geometry guidance can be added, and such over-fitting can be well-constrained. In this way, the correct scene appearance can be learned. The structure of explicit MSI volumes and implicit NeRF rendering also works as a perfect geometry proxy, keeping the network away from over-fitting.
IV Experiments
IV-A Implementation Details
We implement our method using the Pytorch framework. Training and validation are all carried out on a GPU server with four RTX 2080Ti. During optimization, we use ADAM [24] optimizer with a learning rate of $3\times 10^{-4}$. The loss weight of depth is set to $\lambda_{d}=5$. We first train the base model on a generic dataset for $30$ epochs and then fine-tune it on another specific dataset for $10$. We set the inverse depth layer number to $64$ and the range of layer value to $\left[0,2\right]m^{-1}$. Due to limited GPU memory, after inference of geometry and appearance volumes, we randomly sample $8192$ rays on fisheyes and $8192$ rays on the omnidirectional depth map, expressed as $|\mathcal{R}_{p}|=|\mathcal{R}_{f}|=8192$. Then, volume rendering is done on these rays for further color and depth supervision.
For comparison, we keep the network capacity and hyperparameters of baseline methods consistent with our implementation. The same training strategy is used to match the data format. In particular, we add a depth regressor to the output module of MatryODShka [18] so that it can be trained using depth data.
IV-B Datasets
We use the fisheye panorama datasets proposed in OmniMVS [3] for network training. Specifically, we use OmniThings for base model training and OmniHouse for indoor scene fine-tuning. In addition to the four surrounding fisheye cameras, each sample contains a depth ground-truth image in the center. The last $5\%$ of the dataset is extracted for validation.
We also capture fisheye and multi-view images in the Replica [25] simulator to verify the novel view synthesis performance. The Replica360 dataset we constructed contains $33$ indoor scenes, each with $4$ fisheye source views and $5$ randomly sampled omnidirectional target views. The visualization of our Replica360 dataset is in Fig. 3.
Figure 3: Visualization of our Replica360 dataset. The fisheye source views (a), and the multi-view omnidirectional target images (b) are captured from the Replica simulator.
Figure 4: Qualitative results of our depth estimation experiment. We compare the generated omnidirectional depth map with the ground truth depth from the dataset. Our method can generate fine-grained depth estimations while getting rid of over-fitting due to rich texture.
TABLE I: Quantitative evaluation of omnidirectional depth estimation.
Dataset OmniHouse OmniThings
Method $>$ 0.1 $>$ 0.3 $>$ 0.5 MAE RMSE $>$ 0.1 $>$ 0.3 $>$ 0.5 MAE RMSE
MatryODShka 2.091 0.271 0.066 0.017 0.032 12.989 4.046 1.662 0.053 0.032
OmniMVS 1.253 0.117 0.071 0.011 0.023 10.592 3.265 1.305 0.046 0.105
Ours 1.255 0.180 0.033 0.011 0.024 10.149 3.212 1.328 0.044 0.103
Figure 5: Qualitative results of our novel view synthesis experiment. We generate novel panoramic images in the dataset’s target view location and compare them with the ground truth. Our method can generate high-quality and consistent rendering results, avoiding blurring and ghosting artifacts.
IV-C Depth Evaluation
We evaluate the omnidirectional depth estimation performance on a validation set extracted from the OmniHouse and OmniThings. Original and fine-tuned models are used for different datasets, respectively. The qualitative and quantitative evaluation results are shown in Fig. 4 and Tab. I. We use inverse depth value as our comparison metric, and the unit is set to $m^{-1}$. The qualifier ”$>k$” represents the pixel ratio whose error is larger than threshold $k$.
It can be seen that the depth map generated by MatryODShka [18] has blurring effect and poor accuracy. Although OmniMVS [3] can generate fine object edges, it focuses too much on texture, resulting in incorrect depth estimation on surfaces with rich textures. Our method is able to generate fine edges while maintaining overall accuracy.
IV-D Novel View Synthesis Evaluation
To test the novel view synthesis ability, we conduct comparative experiments against MatryODShka [18] and NeRF-360-NGP [6, 16, 20]. For MatryODShka, we replace the original omni-stereo with fisheye images to align the input format. For NeRF-360-NGP, we expand the NeRF method to $360^{\circ}$ using spatial parameterization from Mip-NeRF 360 [16] and achieve efficient training through Instant-NGP [20].
We show our qualitative and quantitative results in Fig. 5 and Tab. II. It can be seen that the color blending design in MatryODShka generates ghosting effect when trained with only depth ground truth. NeRF-360-NGP fails to generate reasonable novel views with limited input images and shared viewing angles. Our method exceeds other generalizable and scene-specific methods in terms of image quality. It achieves high accuracy on novel view synthesis and shows excellent generalization ability on cross-dataset validation.
TABLE II: Quantitative evaluation of panoramic novel view synthesis.
Method Generalizable PSNR SSIM LPIPS
MatryODShka 29.665 0.584 0.419
NeRF-360-NGP 28.594 0.451 0.541
Ours 37.302 0.957 0.077
IV-E Ablation Studies
Our generalizable neural rendering method does not follow traditional novel view synthesis process. It is mainly reflected in three key designs: additional color projection, radiance field rendering, and separated geometry and appearance decoder. We conduct ablation studies to demonstrate the effect of these designs by removing them individually. The results are shown in Fig. 7 and Tab. III.
Although generating good depth maps, the method without additional color input struggles at color learning, producing blurry images. Method without radiance field rendering directly outputs color and occupancy spheres, causing ghosting effects in the synthesized views. Method without separated decoders performs well at novel view synthesis tasks but lacks the ability to estimate fine-grained depth maps. These results show that both designs play an essential role in our method. Ours full model uses color projection input to gain additional color information and implicitly mix them to achieve more reasonable color output. The usage of radiance field rendering creates a more consistent image generation and results in finer color and depth generation. Moreover, the design of separated geometry and appearance decoder makes the role of the network clearer and avoids data confusion. These ablation experiments prove the rationality and innovation of our framework.
IV-F Real-World Experiment
Figure 6: (a): Our self-made multi-fisheye camera rig system for real-world data capture. It comprises $4$ USB fisheye cameras with $220^{\circ}$ FoV arranged on a circle with a diameter of $0.4m$. (b): The captured source views, the generated omnidirectional depth map, and the synthesized image in the real-world experiment.
To test our approach’s generalizability and real-world performance, we constructed a self-collected dataset using a multi-fisheye camera rig with similar settings to the synthetic datasets. The data capture system, fisheye image samples, and the generated results can be seen in Fig. 6. From the generated depth and novel view images, we can see that our method can be well generalized from sim to real. The real-world experiment also demonstrates the practical application value of this method for virtual reality video shooting and 3D scene reconstruction in real-world scenes.
V Conclusion
This paper explores the problem of generating a panoramic scene representation using only four fisheye images as input. To solve this, our presented MSI-NeRF combines the advantage of multi-sphere image representation with volume rendering to construct a generalizable NeRF. From such representation, accurate omnidirectional depth estimation and high-quality novel view synthesis can be conducted. Our proposed semi-self-supervision strategy allows the network to be trained using only depth ground truth. Furthermore, generalizability also enables it to reconstruct unknown scenes efficiently, which is of great significance for robotics and virtual reality applications. As a result, it outperforms baseline methods in both image quality and feasibility.
Figure 7: Qualitative results of our ablation study. We generate novel view images and depth maps of our method without key designs and compare them with our full model. Ours full model generates the finest novel view images while maintaining the depth map quality.
TABLE III: Quantitative evaluation of ablation study. We compare both omnidirectional depth and novel view synthesis metrics.
Method $>$ 0.1 $>$ 0.3 $>$ 0.5 MAE RMSE PSNR SSIM LPIPS
w/o Color Input 1.021 0.140 0.030 0.012 0.024 32.080 0.783 0.287
w/o Radiance Field 1.152 0.308 0.145 0.012 0.023 29.620 0.553 0.432
w/o Separated Decoder 1.396 0.266 0.079 0.013 0.026 37.200 0.953 0.088
Ours Full Model 1.255 0.180 0.033 0.011 0.024 37.302 0.957 0.077